OBSERVATION OF MULTIPLE PROTON/HYDROGEN ATOM RELAY IN MICROSOLVATED 7-AZAINDOLE CLUSTERS IN THE GAS PHASE

\underline{KENJI SAKOTA}, NAOMI INOUE, and HIROSHI SEKIYA, Department of Chemistry, Kyushu University, 6-10-1 Hakozaki, Higashi-ku, Fukuoka, Japan.

A excited-state multiple-proton/hydrogen atom relay on the 7-azaindole-methanol clusters (7AI(MeOH)_{n} n=1-3) has been investigated in the gas phase by measuring the mass-resolved vibronic spectra and the dispersed fluorescence spectra of 7AI(MeOH)_{n}. It has been found that only 7AI(MeOH)_{2} cluster shows visible fluorescence due to the tripleproton/hydrogen atom relay in S_{1}. The 7AI(MeOH)_{2} cluster also exhibits the prominent vibrational-mode specific proton/hydrogen atom relay. The excitation of an intermolecular stretching mode with the in-phase cooperative motion of the whole hydrogen-bond network accelerates the proton/hydrogen atom relay.